# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ 3.1 The Net Riemann Integral
(iii) Let $f,g\in L^{1}(\mu)$ with corresponding Riemann sums in the fixed convergent structure $S_{\gamma}(f,\mu),S_{\gamma}(g,\mu)$. As the product is isotone, we have that, for every $\gamma\in\Gamma$ and $i\in I_{\gamma}$,
@@
Therefore, by taking the limit with respect to the Boccuto-Candeloro net convergence structure $\eta_{3}$, and using the second part of axiom $(i)$ of it, it follows that:
@@
(iv) Follows from the same previous techniques and the use of axioms $(i)$ and $(ii)$ of the Boccuto-Candeloro net convergence structure of $Z$.
We enounce the following direct consequence of Definition 3.9 and Proposition 2.1, which will be useful in more specialized discussions in the next sections:
Corollary 3.1.1.
Suppose $(Z,\eta_{3})$ is a Dedekind Complete Riesz space with $\eta_{3}$ being given by order convergence. Then the integral of a function $f:\Omega\rightarrow X$ with respect to a set function $\mu:\mathcal{H}\rightarrow Y$ exists with respect to $\eta_{3}$ if and only if:
@@
    The net $\{S_{\gamma}(f,\mu)\}_{\gamma\in\Gamma}$ is bounded in $Z$ and,
@@
Having proved these basic consequences of the Net Riemann integral, we offer some more clarifying comments on the integral and its properties.
Observation 3.2.
In this generality, some "expected" properties of the Net Riemann integral may not be valid, such as:
@@
    In general, the Riemann sums $S_{\gamma}(f,\mu)$ will not be monotone, even if $f$ and $\mu$ are positive, in such a way that the integral above cannot be characterized as an infimum/supremum of these nets.
@@
    Considering classical, even Lebesgue, integration theory, a common property of the integrals is given by integration of a indicator function:
    $\int_{\Omega}\mathbbm{1}_{A}d\mu=\mu(A),$
    for some notion of integration. In general, for the Net Riemann integral, such relation is not valid. For its validity, more properties have to be asked on $\mu$, which in general are highly restrictive (e.g, supposing that $(\Omega,\mathcal{H})$ is a topological measure space and $\mu$ is outer/inner regular), or are valid only on some specific cases of such general integration procedure. This will be important later.
@@
    The continuity properties of such Net Riemann integrals are even more delicate, as the integral is given in a non-Lebesgue (i.e, as a net of integrals of an approximating sequence of (simple, bounded,etc) functions) way, making difficult discussions on the passing of limits inside of the integration sign, something already present is definitions of Riemann integration of Henstock-Kurzweil, $S^{*}$, Sion and others.
  4. (iv)
    The finiteness restriction of the nets defining the integration procedure can be restrictive in some cases, making the definition not amenable to procedures such as the $S^{*}$-integral of [14] which deals with countable sums based on nets of partitions and related points in it.
Having point $(iv)$ of the previous observations in mind, we pass now to a brief discussion of infinite sums in the context of Riesz spaces with Boccuto-Candeloro net convergence structures, the $\sigma$-additivity of positive set functions/measures with values in such spaces, and the remaining definitions of integration procedures used in this work. After these definitions, we provide a possible solution to problem (iv) in relation to the Sion [60] integral in a specific type of convergence.
3.2 Summability in Riesz Spaces and a generalized Kolmogorov Integral
We basically have two notions of converging series, one corresponding to unconditional convergence, and other to conditional convergence, as in the classical reference of Bourbaki [46]. In this context, consider the next definition:
Definition 3.10 (Unconditionally Convergent Series).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$. Then, we say that a net $\{x_{i}\}_{i\in I}$ in $X$, with partial (unordered) sums of elements (in $X$)
$\{\sum_{j\in J}x_{j}\}_{J\in\mathcal{F}(I)},$
$x=\lim_{(\mathcal{F}(I),\subseteq,\eta)}\sum_{j\in J}x_{j}$
Definition 3.11 (Conditionally Convergent Series).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$. Then, we say that a sequence of elements $\{x_{n}\}_{n\in\mathbb{N}}$, with partial (ordered) sums of elements (in $X$)
$\{\sum_{i=1}^{n}x_{i}\},$
$x=\lim_{(\mathbb{N},\eta)}\sum_{i=1}^{n}x_{i}$
Similar notions, specifically the first one, can be found in [28], [61] and [62]. By using a cofinal map (see [28]), or a quasi-subnet argument in the terminology of [27], we may prove the following two results:
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$ and $\{x_{n}\}_{n\in\mathbb{N}}$ a sequence with unconditionally convergent series. Then, it has a conditionally convergent series with same limit.
Regarding to the tail of a series, we have:
Proposition 3.6 (Tail of Series).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$ and $\{x_{n}\}_{n\in\mathbb{N}}$ a sequence with a conditionally convergent series. Then, the tail of the series, i.e., the sequence:
$\{\sum_{i=n}^{\infty}x_{i}\}_{n\in\mathbb{N}},$
is convergent (with respect to $\eta$) to 0.
Both results, specially Proposition 3.5, are interesting in the context of measure theory in partially ordered spaces with convergence structures. Regarding this context, the next definition is also useful:
Definition 3.12 (Monotonically Convergence Property).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$. Then, we say that $(X,\eta)$ has the monotonically convergence property (MCP) if, given an increasing bounded above net with a supremum $\{x_{i}\}_{i\in I}$, it converges to its supremum in $\eta$, i.e.
$\lim_{(I,\eta)}\{x_{i}\}_{i\in I}=\sup_{i\in I}x_{i}$
We have the following classical result, which can be found in various references, such as [30], [50], [36], [12], [11] (for further examples, see specially [12]):
Proposition 3.7 (Spaces with the MCP).
If $X$ is a Dedekind complete Riesz space/lattice ordered group with Boccuto-Candeloro convergence structure $\eta$ being $(o_{3})$ or $(o)$ convergence, then it has the MCP property.
We will not provide a proof, as it is standard and can be found in the above references.
We also have the following example of a topological convergence with the MCP:
Example 3.1.
Let $(X,\eta)$ be the space $L^{0}(\Omega,\Sigma,\mathbb{P})$ of (equivalence classes of) real Borel-measurable functions defined on a probability space, equipped with convergence in probability, and the partial ordering of almost sure pointwise inequality. It is not difficult to check that this convergence is topological and gives a Boccuto-Candeloro net convergence structure in $X=L^{0}(\Omega,\Sigma,\mathbb{P})$. In this case, taking $\{X_{n}\}_{n\in\mathbb{N}}$ a monotone sequence of random variables in $X$ bounded above (by a random variable $Y$), it follows from Theorem 3.5 of [63] that this sequence has a limit in probability given by its supremum. It follows that the present $(X,\eta)$ has the MCP.
In this case, as we will use repeatedly, we have the following result, which can be proved as in Lemma 1.2 of [30]:
Proposition 3.8 (Non-negative Series Convergence).
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$ with MCP property and $\{x_{n}\}_{n\in\mathbb{N}}$ a non-negative sequence with conditionally convergent series and partial sums $\{\sum_{i=1}^{n}x_{i}\}_{n\in\mathbb{N}}$ bounded above in $X$. Then, the series associated to $\{x_{n}\}_{n\in\mathbb{N}}$ is unconditionally convergent and:
$\sum_{j=1}^{\infty}x_{i}=\lim_{(\mathcal{F}(\mathbb{N}),\subseteq,\eta)}\sum_{j\in J}x_{j}=\sup_{n\in\mathbb{N}}\sum_{j=1}^{n}x_{i}$
Observation 3.3 (Ideal Infinite Element).
We can dispense with the existence of the supremum, or the MCP property, in the results above by annexing an ideal element $\infty$ to $X$ in such a way that every increasing sequence has a supremum equal to it.
Proposition 3.9.
Let $(X,\eta)$ be a Dedekind complete Riesz space where $\eta$ is the Boccuto-Candeloro net convergence structure given by $(D)$-convergence. Then, a non-negative sequence $\{x_{n}\}_{n\in\mathbb{N}}$ has a conditionally convergent series if and only if it has an unconditionally convergent series with the same limit.

### v2

@@ Abstract
(ii) Let $f,g\in L_{NR}(\mu)$ with corresponding Riemann sums in the fixed convergent structure $S_{\gamma}(f,\mu),S_{\gamma}(g,\mu)$. As the product is isotone, we have that, for every $\gamma\in\Gamma$ and $i\in I_{\gamma}$,
@@
Therefore, by taking the limit with respect to the partially ordered net convergence structure $\eta_{3}$, and using that it is locally solid, it follows that:
@@
(iii) Follows from the same previous techniques and the use of axioms $(i)$ and $(ii)$ of the partially ordered net convergence structure of $Z$.
These results give the basic properties of the abstract integral defined in this section that are valid in full generality. We now aim to prove some specific results concerning hereditary (i.e subset) integration, that is: if a function $f$ is Net Riemann integrable on $B\in\mathcal{H}$ and $A\in\mathcal{H}$ is such that $A\subseteq B$, does it implies that $f$ is integrable on $A$ ? As the example of the Riemann-Stieltjes integral of McShane (1969) shows, this cannot be positively answered in general. To obtain such result, we impose further conditions on the Szábo-Száz net that gives the integration procedure (which are satisfied for general integration procedures such as McShane (1969) and Sion (1973)).
Theorem 2.4 (Subset Integration).
Suppose that $Z$ is complete in its net convergence structure. Fix $A,B\in\mathcal{H}$ such that $A\subseteq B$ and suppose that either $\mathcal{H}$ is closed under taking complements of its members or that $A^{c}\in\mathcal{H}$ . Consider the following hypotheses:
@@
    For each $\gamma_{B},\gamma_{B}^{{}^{\prime}}\in\Gamma(B)$, there exists $\gamma_{A},\gamma^{{}^{\prime}}_{A}\in\Gamma(A)$ such that, for each $\gamma^{1}_{A},\gamma^{2}_{A}\in\Gamma(A)$ that satisfies $\gamma^{1}_{A}\geq\gamma_{A},\gamma^{2}_{A}\geq\gamma^{{}^{\prime}}_{A}$, we can find $\gamma_{A^{c}}\in\Gamma(A^{c})$ in such a way that there exists $\gamma^{1}_{B},\gamma^{2}_{B}\in\Gamma(B)$, such that $\gamma^{1}_{B}\geq\gamma_{B},\gamma^{2}_{B}\geq\gamma_{B}$, with corresponding Fleischer-Szabó-Száz nets
    $\gamma^{1}_{B}\rightarrow((\sigma_{\gamma^{1}_{A}}\cup\sigma_{\gamma_{A^{c}}}),(\tau_{\gamma^{1}_{A}}\cup\tau_{\gamma_{A^{c}}})),$
    $\gamma^{2}_{B}\rightarrow((\sigma_{\gamma^{2}_{A}}\cup\sigma_{\gamma_{A^{c}}}),(\tau_{\gamma^{2}_{A}}\cup\tau_{\gamma_{A^{c}}})),$
@@
    For each $\gamma_{B}\in\Gamma(B)$ there exists $\gamma_{A}\in\Gamma(A)$ and $\gamma_{A^{c}}\in\Gamma(A^{c})$ such that for each $\gamma^{{}^{\prime}}\in\Gamma(A)$ and $\gamma^{{}^{\prime\prime}}\in\Gamma(A^{c})$, there exists $\gamma\in\Gamma(B)$ of the form
    $\gamma_{B}\rightarrow((\sigma_{\gamma^{{}^{\prime}}}\cup\sigma_{\gamma_{{}^{\prime\prime}}}),(\tau_{\gamma^{{}^{\prime}}}\cup\tau_{\gamma_{{}^{\prime\prime}}})),$
    and $\gamma\geq\gamma_{B}$.
Then, if $f$ is net Riemann integrable on $B$, it is net Riemann Integrable on $A$ and $A^{c}$ (the complement relative to $B$) and:
$\int_{B}fd\mu=\int_{A}fd\mu+\int_{A^{c}}fd\mu$
Proof.
As $f$ is net Riemann integrable on $B$, the net of Riemann sums differences formed on $\Gamma(B)\times\Gamma(B)$, with values in $Z$,
$\{\sum_{i\in I_{\gamma^{{}^{\prime}}}}f(\tau_{\gamma^{{}^{\prime}}_{i}})\mu(\sigma_{\gamma^{{}^{\prime}}_{i}})-\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})\}_{(\gamma,\gamma^{{}^{\prime}})\in\Gamma(B)\times\Gamma(B)},$
is Cauchy in the net convergence structure of $Z$. Fix $\gamma_{B},\gamma^{{}^{\prime}}_{B}\in\Gamma(B)$. Now, by the first hypothesis (i), there exists $\gamma_{A},\gamma^{{}^{\prime}}_{A}\in\Gamma(A)$ such that, for each $\gamma^{1}_{A},\gamma^{2}_{A}\in\Gamma(A)$ that satisfies $\gamma^{1}_{A}\geq\gamma_{A},\gamma^{2}_{A}\geq\gamma^{{}^{\prime}}_{A}$, we can find $\gamma_{A^{c}}\in\Gamma(A^{c})$ in such a way that there exists $\gamma^{1}_{B},\gamma^{2}_{B}\in\Gamma(B)$, such that $\gamma^{1}_{B}\geq\gamma_{B},\gamma^{2}_{B}\geq\gamma_{B}$, with corresponding Fleischer-Szabó-Száz nets
$\gamma^{1}_{B}\rightarrow((\sigma_{\gamma^{1}_{A}}\cup\sigma_{\gamma_{A^{c}}}),(\tau_{\gamma^{1}_{A}}\cup\tau_{\gamma_{A^{c}}})),$
$\gamma^{2}_{B}\rightarrow((\sigma_{\gamma^{2}_{A}}\cup\sigma_{\gamma_{A^{c}}}),(\tau_{\gamma^{2}_{A}}\cup\tau_{\gamma_{A^{c}}})),$
Which gives the following decomposition of Riemann sums:
$\sum_{i\in I_{\gamma^{1}_{B}}}f(\tau_{\gamma^{1,B}_{i}})\mu(\sigma_{\gamma^{1,B}_{i}})=\sum_{j\in I_{\gamma^{1}_{A}}}f(\tau_{\gamma^{1,A}_{j}})\mu(\sigma_{\gamma^{1,A}_{j}})+\sum_{k\in I_{\gamma_{A^{c}}}}f(\tau_{\gamma^{A^{c}}_{k}})\mu(\sigma_{\gamma^{A^{c}}_{k}}),$
$\sum_{i\in I_{\gamma^{2}_{B}}}f(\tau_{\gamma^{2,B}_{i}})\mu(\sigma_{\gamma^{2,B}_{i}})=\sum_{j\in I_{\gamma^{2}_{A}}}f(\tau_{\gamma^{2,A}_{j}})\mu(\sigma_{\gamma^{2,A}_{j}})+\sum_{k\in I_{\gamma_{A^{c}}}}f(\tau_{\gamma^{A^{c}}_{k}})\mu(\sigma_{\gamma^{A^{c}}_{k}}).$
Therefore, $\{\sum_{i\in I_{\gamma^{1}_{A}}}f(\tau_{\gamma^{1,A}_{i}})\mu(\sigma_{\gamma^{1,A}_{i}})-\sum_{i\in I_{\gamma^{2}_{A}}}f(\tau_{\gamma^{2,A}_{i}})\mu(\sigma_{\gamma^{2,A}_{i}})\}_{(\gamma^{1}_{A},\gamma^{2}_{A})\geq(\gamma_{A},\gamma^{{}^{\prime}}_{A})\in\Gamma(A)\times\Gamma(A)}$ is contained in $\{\sum_{i\in I_{\gamma_{B}^{1}}}f(\tau_{\gamma^{1,B}_{i}})\mu(\sigma_{\gamma^{1,B}_{i}})-\sum_{i\in I_{\gamma^{2}_{B}}}f(\tau_{\gamma^{2,B}_{i}})\mu(\sigma_{\gamma^{2,B}_{i}})\}_{(\gamma^{*}_{B},\gamma^{**}_{B})\geq(\gamma_{B},\gamma^{{}^{\prime}}_{B})\in\Gamma(B)\times\Gamma(B)}$, which shows that the net of differences of Riemann sums in $\Gamma(A)\times\Gamma(A)$ is a quasi-subnet of net of differences of Riemann sums in $\Gamma(B)\times\Gamma(B)$ . By the first paragraph (completeness of the convergence structure of $Z$) and axiom (ii) of Definition 1.13, we conclude that $f$ is integrable on $A$. A similar argument gives that $f$ is integrable on $A^{c}$.
Similarly, by (ii), it follows $\{\sum_{j\in I_{\gamma_{A}}}f(\tau_{\gamma^{A}_{j}})\mu(\sigma_{\gamma^{A}_{j}})+\sum_{k\in I_{\gamma_{A^{c}}}}f(\tau_{\gamma^{A^{c}}_{k}})\mu(\sigma_{\gamma^{A^{c}}_{k}})\}_{(\gamma_{A},\gamma_{A^{c}})\in\Gamma(A)\times\Gamma(A^{c})}$ is a quasi-subnet of $\{\sum_{j\in I_{\gamma_{B}}}f(\tau_{\gamma^{B}_{j}})\mu(\sigma_{\gamma^{B}_{j}})\}_{\gamma_{B}\in\Gamma(B)}$, in such a way that by axiom (ii) of Definition 1.13, they convergence to the same limit, $\int_{B}fd\mu$. Now, by the double indexing technique used in (O’Brien et al., 2023, Remark 2.5, p. 7), we may regard the Riemann sums of $f$ in $A$ and $A^{c}$ as convergent nets indexed in $\Gamma(A)\times\Gamma(A^{c})$, converging to $\int_{A}fd\mu$ and $\int_{A^{c}}fd\mu$ respectively. Combining this results, we have:
$\int_{B}fd\mu=\lim_{(\Gamma(A)\times\Gamma(A^{c}),\eta_{3})}(\sum_{j\in I_{\gamma_{A}}}f(\tau_{\gamma^{A}_{j}})\mu(\sigma_{\gamma^{A}_{j}})+\sum_{k\in I_{\gamma_{A^{c}}}}f(\tau_{\gamma^{A^{c}}_{k}})\mu(\sigma_{\gamma^{A^{c}}_{k}}))$
$\lim_{(\Gamma(A)\times\Gamma(A^{c}),\eta_{3})}\sum_{j\in I_{\gamma_{A}}}f(\tau_{\gamma^{A}_{j}})\mu(\sigma_{\gamma^{A}_{j}})+\lim_{(\Gamma(A)\times\Gamma(A^{c}),\eta_{3})}\sum_{k\in I_{\gamma_{A^{c}}}}f(\tau_{\gamma^{A^{c}}_{k}})\mu(\sigma_{\gamma^{A^{c}}_{k}})=\int_{A}fd\mu+\int_{A^{c}}fd\mu$
Therefore, from this last result, we get the following additivity property of the net Riemann integral:
Corollary 2.4.1 (Additivity of the Integral).
Let $f$ be an element of $L_{NR}(\mu)$. Then, under the hypotheses of Theorem 2.4, the set function:
$A\in\mathcal{H}\rightarrow\eta_{f}(A)=\int_{A}fd\mu,$
with values in $Z$, is an (finitely) additive set function defined on $\mathcal{H}$.
Specific applications of these results in the topological case can be found in the integrals of (Sion, 1973, Theorem 4.2, p. 18) and Millington (2005), and in the partially ordered case (in the Henstock-Kurzweil case) of (Riečan and Neubrunn, 1997, Proposition 5.2.10, p. 79). One other possibility to obtain the same result is to ask, as did Fleischer (1993), that the directed sets $\Gamma(A\cup B)$ and $\Gamma(A)$ (and $\Gamma(B)$), for $A,B$ disjoint, are isomorphic, in the sense of always being possible to obtain a partition from the union of disjoint sets to each one individually, and vice versa. The previously cited integrals satisfy this property, and it implies the conditions in Theorem 2.4. Therefore, our result is more general.
Corollary 2.4.1 notwithstanding, the behavior of the set function induced by a net Riemann integrable functions is not necessarily simple. An interesting property would be that the function $f=\mathbbm{1}_{A}$, for $A\in\mathcal{H}$, is integrable with respect to $\mu$ and:
$\nu_{f}(A)=\mu(A).$
In general, that is not true for integrals such as the Henstock-Kurzweil integral for vector-valued (specially operator-valued) measures (see the discussion in Chapter 5 of Riečan and Neubrunn (1997)), except under regularity conditions of the Radon type. As our integral contains this one as a special case (which well be shown shortly), the net Riemann integral will not have this property. Having this problem in mind, Szabó and Száz (1988) isolated the integration of indicator functions as a formal property and defined the next concept:
Definition 2.6 (Regular Integrators).
$\int_{A}\mathbbm{1}_{\Omega}d\mu=\mu(A),$
and,
$\int_{\Omega}\mathbbm{1}_{A}d\mu=\mu(A)$
for each $A\in\mathcal{H}$.
It follows directly from the linearity properties of the net Riemann integral in Corollary 2.4.1 that simple functions (to be defined next) are then net Riemann Integrable and the usual additivity property holds, that is:
Proposition 2.1 (Integration of Simple Functions).
Let $f:\Omega\rightarrow X$ be a $\mathcal{H}$-simple function, i.e:
$f=\sum_{i=1}^{n}x_{i}\mathbbm{1}_{A_{i}},$
$\int_{\Omega}fd\mu=\sum_{i=1}^{n}x_{i}\mu(A_{i})$
The next strengthening of these two last properties, hereditary/additivity of integration and regular integrators, would be to prove that, if $\mu$ is $\sigma$-additive in some appropriate sense, then the set function induced by a net Riemann integrable function $f$, given by $\nu_{f}$, is $\sigma$-additive in $\mathcal{H}$. As the case of the Henstock-Kurzweil integral of a Banach-valued function with respect to a scalar $\sigma$-additive measure shows (see (Schwabik and Ye, 2005, Proposition 5.4.1, p. 159) or Swartz (2004)), that is not true in general. This is related to the fact that we may construct integrals by the procedure of Definition 2.4 that includes refinements and gauge integrals using finite partitions, and therefore based only on strictly finite Riemann sums, and that are also of non-absolute type. That is in contrast with integrals such as Rickart (1942), Kolmogoroff (1930) and Sion (1973) integrals, which are based on the refinement and truncation order in countable partitions/coverings of the base space, having the $\sigma$-additivity of the integral-induced measure naturally (a consequence of the unconditional summability of the approximating sums involved in the limit process, inducing an absolute integral). As we will see in the section about a generalization of Kolmogoroff (1930) $S^{*}$ integral to net convergence structures, obtaining a property of this type ($\sigma$-additive measure implies $\sigma$-additive integral-induced measure) is possible even without imposing a topology or order in the image space $Z$.
We now pass to a brief discussion of the properties of the space $L_{NR}(\mu)$. An immediate consequence of Corollary 2.4.1 and Proposition 2.1 is the vector space structure of this space:
Proposition 2.2 ($L_{NR}(\mu)$ is a Vector Space).
Let $L_{NR}(\mu)$ be the space of net Riemann integrable functions with respect to $\mu$ in a fixed integration structure. Then, $(L_{NR}(\mu),+)$ is a semigroup additively written with respect to the ++ operation of $X$, and $(L_{NR}(\mu),+,(\mathbb{R},\cdot))$ is a vector space with pointwise product by real numbers  [^5] As $X$ is a vector space and the bilinear product is compatible with pointwise multiplication in $X$ by real numbers.. If $\mu$ is a regular integrator, $L_{NR}(\mu)$ contains the space of $X$-valued simple functions as a vector subspace.
In the case of $X,Y,Z$ have a fixed partially ordered net convergence structure, $L_{RM}(\mu)$ is a partially ordered vector space with the pointwise order for functions: that is, $f,g\in L_{RM}(\mu)$ are such that $f\leq g$ if and only if $f(\omega)\leq g(\omega)$ (in $X$) for each $\omega\in\Omega$.
At this level of generality, not much more interesting can be said about this space. For example, we were unable to prove continuity properties of the integration functional for $L_{NR}(\mu)$ without imposing some more conditions, not necessarily of topological or order-convergence properties, on the convergence structures of $X,Y$ and $Z$ (see the next sections). The action of (real) linear functions on $L_{NR}(\mu)$, related to a type of Pettis integral, can be elucidated by some analogue constructions as in topological/partially ordered vector spaces, but we will comment about this aspect in the last part of the present section.
In the further case of $X,Y,Z$ on a fixed partially ordered net convergence structure, two interesting questions should be asked about $L_{NR}(\mu)$:
  1. 1.
    Is it, or can it be made, a Lattice ?
  2. 2.
    In general, if $f:\Omega\rightarrow X$ is a function such that $|f|\leq|g|$ in the pointwise partial order (as in the definition above), and $g\in L_{NR}(\mu)$, does it implies that $f\in L_{RM}(\mu)$ ? In other words, is $L_{NR}(\mu)$ solid in the space of all functions of the type $f:\Omega\rightarrow X$ ?
As is expected, the Henstock-Kurzweil integral provides a counterexample to these claims for the net Riemann integral. As this integral is non-absolute, if $X=Z$ a Banach lattice and $Y=\mathbb{R}_{+}$ and $\mu$ the Lebesgue measure, its space of integrable functions, equipped with the Alexiewicz norm, is not solid in the sense above (see, for example, Federson (2004)). Also, by the same reason, it is not a lattice. Some results in the lattice direction seem possible by using a modification of the lattice-integration results of Hayes (1974) to the present case, but we leave the details for further work.
We now discuss the Pettis integral aspect of the net Riemann integral, the last property in the general context of net convergence structures. For this, we mainly use the material in Chapters 3-5 of Beattie and Butzmann (2002), but adapted to nets. Related material (in the filter language) can be found in Lucyshyn-Wright (2012) and Lucyshyn-Wright (2013).
We begin with the following definition:
Definition 2.7 (Dual of a Net Convergence Space).
Let $(X,\eta)$ be a topological vector space with a net convergence structure $\eta$. Then, the space of all linear continuous real valued functionals $\psi:X\rightarrow\mathbb{R}$, where $\mathbb{R}$ is equipped with its canonical (euclidean) topology, is called the continuous dual of $X$, which we denote by $\mathcal{L}(X)$.
We assume, tacitly, that $\mathcal{L}(X)$ is non-empty in each statement involving it. Its now easy to see that:
Proposition 2.3 (Continuous Convergence).
Let $\mathcal{L}(X)$ be the continuous dual of the net convergence space $(X,\eta)$, where $X$ is a topological vector space. Consider the next definition of convergence, which we denote by $\eta_{\mathcal{L}(X)}$: a net $\{\psi_{i}\}_{i\in I}$ of elements of $\mathcal{L}(X)$ converges to $\psi\in\mathcal{L}(X)$ if and only if for every $x\in X$ fixed, $\{\psi_{i}(x)\}_{i\in I}$ converges to $\psi(x)$. Then, $(\mathcal{L}(X),\eta_{\mathcal{L}(X)})$ is a net convergence space.
For the rest of this section, we assume that the continuous dual of every net convergence space is equipped with the convergence of the last proposition.
For the definition of the Pettis integral in this context, we shall need the next condition (see (Beattie and Butzmann, 2002, Lemma 4.2.5, p. 127)):
Definition 2.8 (Separated Continuous Dual).
Let $\mathcal{L}(X)$ be the continuous dual of the net convergence space $(X,\eta)$, where $X$ is a topological vector space. We say that $\mathcal{L}(X)$ separates points of $X$ if:
$\bigcap\{ker(\psi):\psi\in\mathcal{L}(X)\}=\{0\}$
By (Beattie and Butzmann, 2002, Lemma 4.2.5, p. 127), this is equivalent to ask that the canonical embedding of $X$ into its (continuous convergence) double dual is injective. For now on in this section, we shall assume that each $\mathcal{L}(X)$ is separated.
The last convergence-theoretic notion is the following:
Definition 2.9 (Weak Convergence in $X$).
Let $\mathcal{L}(X)$ be the continuous dual of the net convergence space $(X,\eta)$, where $X$ is a topological vector space. Then, we say that $X$ is equipped with the weak net convergence structure $\eta_{w}$ when a net $\{x_{i}\}_{i\in I}$ converges to $x\in X$ if and only if for each $\psi\in\mathcal{L}(X)$, the (real valued) net $\{\psi(x_{i})\}_{i\in I}$ converges to $\psi(x)$ in the euclidean topology of $\mathbb{R}$. We denote this net convergence space by $(X,\eta_{w})$.
Suppose now that $(\Omega,\mathcal{H},\mu)$ is a fixed complete finite measure space, and that $(X,\eta)$ is a net convergence space, where $X$ is a topological vector space. As said previously, we shall also suppose that $\mathcal{L}(X)$ is non-empty [^6] This is not automatic, as the case of non-locally convex topological vector spaces show. See Adasch et al. (1978). and separated.
We now have the next essential definition:
Definition 2.10 (Weakly Measurable Function).
We say that a function $f:\Omega\rightarrow X$ is weakly measurable if for every $\psi\in\mathcal{L}(X)$, the function $\psi(f):\Omega\rightarrow\mathbb{R}$, defined by $\psi(f)(\omega)=\psi(f(\omega))$ for each $\omega\in\Omega$, is a Borel measurable function.
We also have, associated to this last definition:
Definition 2.11 (Scalarly Integrable Function).
We say that a weakly measurable function $f:\Omega\rightarrow X$ is scalarly integrable if:
$\int_{\Omega}^{(L)}\psi(f)d\mu<\infty,$
for each $\psi\in\mathcal{L}(X)$, where this last integral is taken in the Lebesgue sense.
We now state the definition of Pettis mode of integration (for the locally convex topological vector space case, see Jaker Ali and Chakraborty (1997) or Musial (2002)):
Definition 2.12 (Pettis Integral).
Let $f:\Omega\rightarrow X$ be a scalarly integrable function. Then, we say that $f$ if Pettis integrable on $A\in\mathcal{H}$ with Pettis integral $x_{A}$ if there exists an element $x_{A}$ in $X$ such that,
$\psi(x_{A})=\int_{A}^{(L)}\psi(f)d\mu,$
for every $\psi\in\mathcal{L}(X)$ and the integral in the right-hand side taken in Lebesgue sense.
If f is Pettis integrable on every $A\in\mathcal{H}$, then we say that $f$ is Pettis integrable.
In general, we denote the Pettis integral of a map $f$ by:
$\int^{(P)}_{\cdot}fd\mu,$
with the integrating set omitted when no confusion is possible.
We point out now that this notion of "weak integration" contains not only the classical concept of Pettis (1938), but also its partially ordered analogue in Sipos (1981), and the Radon type construction of Lucyshyn-Wright (2012).
To guarantee that the Pettis integral is well defined, we need the next uniqueness result for this integral. As it is straightforward, $\mathcal{L}(X)$ being supposed to be separated, we skip it.
Proposition 2.4 (Uniqueness of the Pettis Integral).
Let $f:S\rightarrow X$ be a Pettis integrable function. Then, its Pettis integral is unique.
We now prove the main result we alluded before, we the cumulative notation fixed until now:
Theorem 2.5 (Net Riemann Integral and Pettis Integration).
Let $f:S\rightarrow X$ be a scalarly integrable function, where $X$ is equipped with its weak convergence structure $\eta_{w}$. Suppose also that if $g:\Omega\rightarrow\mathbb{R}$ is a real valued function (Lebesgue) integrable with respect to $\mu$, then its net Riemann integrable exists coincides with its Lebesgue integral, and conversely [^7] That includes most of the integrals that we shall study. See the next section. Note the hypothesis of weak measurability, which is essential here for this hypothesis too.. Then, we have:
@@
    $f$ is Pettis integrable if and only if $\int_{A}fd\mu\in X$ for each $A\in\mathcal{H}$, where this last integral is the net Riemann integral of $f$;
@@
    f is Pettis integrable on every $A\in\mathcal{H}$ if and only if $f$ is net Riemann integrable;
@@
    We have the identification:
    $\int_{A}^{(P)}fd\mu=\int_{A}fd\mu,$
    for each $A\in\mathcal{H}$.